OpenStudy (anonymous):

About 11% of the population is math phobic. If two people are randomly selected, what is the probability that both are math phobic? What is the probability that at least one is math phobic?

OpenStudy (blues):

P = (0.11)^2 = 0.0121

OpenStudy (anonymous):

first one is \[.11^2\] whatever that is

OpenStudy (anonymous):

at least one means not none. none is \[.89^2\]so answer is \[1-.89^2\]

OpenStudy (anonymous):

First mathfobic AND(multiplication) second mathfobic \[P(A)=\frac{11}{100}\cdot\frac{11}{100}\] To get at leats one mathfobic you can calculate probability of 0 mathfobics and take inverse \[P({B})=1-P(B)=1-\frac{89}{100}\cdot\frac{89}{100}\]

OpenStudy (anonymous):

Thanks! Big help everyone :)

OpenStudy (anonymous):

Sorry I meant \[P(B)=1-P(\bar{B})=...\]